Find the equilibrium le find the equilibrium level of income and interest rate and the levels of C , I , Mt and Mz C= 89+0.6y I =120-150i
1
Expert's answer
2021-10-27T09:16:38-0400

"Y=C+I"

"Y=89+0.6Y+120-150i"

"Y-0.6Y=209-150i"

"0.4Y=209-150i"

Thus, the equilibrium level of income is given by:

"Y=\frac{209-150i}{0.4}" .